The following processes P1, P2, P3, P4 and P5 arrive at the same time (t 0). Establish a timeline of the process scheduling for the following scheduling algorithms while also identifying start times for each process l. FCFs (first come, first serve, assume order of P1, P2, P3, P4 and P5) ll. SJF (shortest job first) Ill. Priority iv Round Robin (quantum F 1, assume order of P1, P2, P3, P4 and P5) Determine the average waiting time for each algorithm Priority Process Burst time/processing time P1 10 P3 P4 P5
Answer #1

1.

FCFS

P1 P2 P3 P4 P5

0---------------------10--------------------11--------------------13------------------14-------------------19

Average Waiting Time = (0+10+11+13+14)/5 = 9.6

2.

SJF

P2 P4 P3 P5 P1

0--------------------1----------------------2---------------------4---------------------9--------------------19

Average waiting time = (9+0+2+1+4)/5 = 3.2

3.

Priority

P2 P5 P1 P3 P4

0--------------------1---------------------6---------------------16---------------------18------------------19

Average waiting time = (6+0+16+18+1)/5 = 8.2

4.

Round Robin(quantum =1)

P1 P2 P3 P4 P5 P1 P3 P5 P1 P5 P1 P5 P1 P5 P1 P1 P1 P1 P1

0----1----2----3----4-----5-----6-----7----8----9----10---11---12---13---14---15---16---17---18--19

Average waiting time = (18+1+6+3+13)/5 = 8.2